Hello,
I am working on word problems in Algebra. Here is on of them.
The amount of time it takes a swimmer to swim a race varies inversely as the average speed of a swimmer. A swimmer finishes a race in 500 seconds with an average speed of three feet per second. Find the Average speed of the swimmer if it takes 375 seconds to finish the race.
How would I do this?!?
Algebra 1

Answer by dolly(163)

Let me explain this to you.
let t be the time in seconds and s be the speed in ft/sec
Given that the amount of time it takes a swimmer to swim a race varies inversely as the average speed of a swimmer.
==> t = K/ s where K is the constant of proportionality. (inverse proportion)
A swimmer finishes a race in 500 seconds with an average speed of three feet per second.
==> 500 = K/3
==> K = 1500 [multiplying by 3 on both the sides]
Thus the original equation becomes, t = 1500/s
Now we need to find the Average speed of the swimmer if it takes 375 seconds to finish the race.
==> 375 = 1500/s
==> 375 s = 1500 [multiplying by s on both the sides]
==> s = 1500/375
==> s = 4
Thus the average speed of the swimmer is 4 ft/sec if he finishes the race in 375 sec.

Best of luck!!!